# CUE-RAG: Towards Accurate and Cost-Efficient Graph-Based RAG via Multi-Partite Graph and Query-Driven Iterative Retrieval

### v1

@@ Abstract.
Despite the remarkable progress of Large Language Models (LLMs), their performance in question answering (QA) remains limited by the lack of domain-specific and up-to-date knowledge. Retrieval-Augmented Generation (RAG) addresses this limitation by incorporating external information, often from graph-structured data. However, existing graph-based RAG methods suffer from poor graph quality due to incomplete extraction and insufficient utilization of query information during retrieval. To overcome these limitations, we propose CUE-RAG, a novel approach that introduces (1) a multi-partite graph index incorporates text Chunks, knowledge Units, and Entities to capture semantic content at multiple levels of granularity, (2) a hybrid extraction strategy that reduces LLM token usage while still producing accurate and disambiguated knowledge units, and (3) Q-Iter, a query-driven iterative retrieval strategy that enhances relevance through semantic search and constrained graph traversal. Experiments on three QA benchmarks show that CUE-RAG significantly outperforms state-of-the-art baselines, achieving up to 99.33% higher Accuracy and 113.51% higher F1 score while reducing indexing costs by 72.58%. Remarkably, CUE-RAG matches or outperforms baselines even without using an LLM for indexing. These results demonstrate the effectiveness and cost-efficiency of CUE-RAG in advancing graph-based RAG systems.
@@ 1. Introduction
Large Language Models (LLMs) like Qwen3 (Yang et al., 2025), DeepSeek (DeepSeek-AI et al., 2025), and LLaMA3.1 (Grattafiori et al., 2024) have received tremendous attention from both industry and academia (Ghimire et al., 2024; Huang and Huang, 2024; Liu et al., 2024; Nie et al., 2024; Wang et al., 2024). Despite their remarkable success in question-answering (QA) tasks, they may still generate wrong answers due to a lack of domain-specific and real-time updated knowledge outside their pre-training corpus (Peng et al., 2024). To enhance the trustworthiness and interpretability of LLMs, Retrieval-Augmented Generation (RAG) methods (Fan et al., 2024; Gao et al., 2023; Hu and Lu, 2024; Huang and Huang, 2024; Wu et al., 2024; Yu et al., 2024; Zhao et al., 2024) have emerged as a core approach, which often retrieve relevant information from documents, relational data, and graph data to facilitate the QA tasks. The state-of-the-art (SOTA) RAG approaches often use the graph data as the external data since they capture the rich semantic information and link relationships between entities.
In the literature, representative graph-based RAG methods typically follow a two-phase pipeline: (1) Offline graph construction: The external text corpora $\mathcal{T}$ are first segmented into smaller chunks, from which nodes and edges are extracted using LLMs to build a knowledge graph (KG) (Zhou et al., 2025; Xiang et al., 2025). (2) Online retrieval: Given an online query $q$, it either retrieves relevant nodes, edges, or subgraphs from the KG for direct answer generation, or optionally traces these elements back to their source text chunks via associated provenance links to support answer generation (Gutiérrez et al., 2024; Edge et al., 2024; Guo et al., 2024). While they offer improved performance, both approaches suffer from two major limitations:
[…]
Limitation 2: Insufficient utilization of the query information. Most existing methods (Huang et al., 2025; Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) for online retrieval follow two steps: (1) use the input query to identify relevant elements (e.g., nodes, edges) in the KG, and (2) retrieve additional information based on the above elements, such as text chunks or subgraphs, to augment the context for answer generation. While the first step ensures query relevance, the second step often fails to fully leverage the query information, resulting in a semantic misalignment between the query and the retrieved context.
@@
Figure 1. Comparison of our extraction method (Below) and rigid triple extraction in existing Graph-based RAG methods (Upper).
@@
To address Limitation 1, we introduce the novel concept of knowledge unit, a statement that conveys an atomic piece of information from a text chunk. Collectively, these knowledge units can fully reconstruct the semantic content of the original text. By extracting entities from the knowledge units, we can effectively address missing nodes or edges in the constructed graph. This leads to the construction of a multi-partite graph composed of three types of nodes, each representing distinct semantic granularities: (1) Chunk, (2) knowledge Unit, and (3) Entity. These nodes are connected through extraction and containment relations, forming a coherent multi-partite graph, as detailed in Section 3.2 and illustrated in Figure 3.
@@
Intuitively, each sentence in a text chunk can serve as a basic knowledge unit, and lightweight NLP tools (Bird et al., 2009) can be used to extract these units. However, this approach overlooks sentence context, leading to contextual ambiguity – semantically similar sentences from different chunks may actually have distinct meanings. For example, as shown in Figure 2, two such sentences extracted by an NLP tool may have high cosine similarity (0.91) despite differing in meaning. To address this, we leverage LLMs to incorporate context when extracting knowledge units. As a result, the LLM-extracted units in the same example show much lower similarity (0.51), effectively capturing context-sensitive distinctions. Clearly, directly using LLMs to extract the units for all text chunks is very time and money-consuming. In this paper, we propose a novel hybrid extraction strategy that combines the advantages of powerful LLMs with lightweight NLP tools (Bird et al., 2009). An ideal strategy would involve an oracle model that can identify which knowledge unit in a chunk may exhibit contextual ambiguity with units in other chunks. When such ambiguity is detected, we can use LLMs to resolve it, while relying on NLP tools for unambiguous units—thereby optimizing both efficiency and accuracy. Although such an oracle model does not exist, we can use semantic similarity between chunks as a proxy to evaluate potential contextual ambiguity, since knowledge units in highly similar chunks are intuitively more prone to ambiguity. To address this, we assign chunks with high semantic similarity to others for LLMs processing, while the remainder use NLP tools. To balance token cost and extraction accuracy, we formulate the chunk assignment as a 0-1 knapsack problem (Sahni, 1975), aiming to maximize the total semantic similarity of the chunks selected for LLM processing within a given token budget. This strategy matches the performance of end-to-end LLM-based unit extraction, but uses only 50% of the tokens (see Experiment 4.3). Remarkably, CUE-RAG matches or outperforms baselines even without using an LLM for indexing (see Experiment 4.2).
Figure 2. Comparison between sentence-level and LLM-extracted knowledge units. While the sentence-level units (top) lack specificity, the LLM-extracted units (bottom) offer more precise and complete semantic representations.
@@
To address Limitation 2, we propose Query-drIven iterative retrieval (Q-Iter), an iterative retrieval strategy that alternately traverses knowledge units and entities. The process begins by extracting entities from the input query to activate related entities, while concurrently using semantic search to anchor associated knowledge units. These units, in turn, expand the entity set by introducing entities referenced in their content. The strategy then performs constrained multihop traversal over the multi-partite graph to retrieve more relevant units. To ensure alignment with the query, each candidate unit is scored using a lightweight re-ranker specialized in query-context relevance. Finally, the collected knowledge units are mapped to their originating text chunks, which are subsequently re-ranked to yield the final context for answer generation.
@@
In our experiments, we evaluate 13 solutions across three datasets, focusing on two key aspects: (1) Cost efficiency, based on token usage during offline indexing and online retrieval; and (2) QA performance, measured by F1 score and Accuracy. Remarkably, CUE-RAG demonstrates significant improvements over the SOTA baselines, delivering up to 99.33% higher Accuracy and 113.51% greater F1 score while simultaneously reducing indexing costs by 72.58%. Remarkably, the zero-token variant of CUE-RAG achieves comparable or superior performance to baselines. This highlights CUE-RAG’s superior graph indexing and inherent retrieval capabilities.
In summary, we make the following contributions through the development and evaluation of CUE-RAG:
@@
  • We propose a multi-partite graph index that integrates text chunks, knowledge units, and entities to support semantically coherent retrieval across multiple levels of granularity.
  • We propose a hybrid strategy for knowledge unit extraction that maximizes LLM processing benefits while minimizing token usage.
@@
  • We empirically validate the effectiveness of CUE-RAG and the proposed retrieval strategy through comprehensive experiments against competitive baselines.
Figure 3. CUE-RAG consists of two phases: offline indexing and online retrieval.
@@ 2. Preliminary
2.1. Terminology
Let $\mathcal{T}$ be a text corpora whose element $t_{i}\in\mathcal{T}$ is a text chunk of token length $\leq l$. The embedding of each text chunk $t_{i}$ is defined as $\bm{\phi}(t_{i})$. We define a graph as $\mathcal{G}=(\mathcal{V,E})$, where $\mathcal{V}$ represents the set of nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of edges connecting pairs of nodes. In addition, we use calligraphic uppercase letters (e.g., $\mathcal{S}$) to represent a set of elements. Table 1 provides a summary of commonly used notations.
@@ 2.1. Terminology
$N$ Returned chunks size $N$
@@
2.2. Background
Representative graph-based RAG methods typically follow two key stages: offline indexing and online retrieval. In the offline phase, a KG $\mathcal{G}$ is constructed from a text corpora $\mathcal{T}$ by extracting salient components, including nodes and edges, together with their associated textual attributes, such as descriptions or keywords (Guo et al., 2024; Edge et al., 2024). These components are then stored in vector databases or graph systems for efficient retrieval (Nebula, 2010; Neo4j, 2006). During online retrieval, the input query $q$ is usually incorporated only at the initial step, using its semantic embedding or extracted entities to match nodes or relations in $\mathcal{G}$ (Gutiérrez et al., 2024; Guo et al., 2024; Huang et al., 2025; He et al., 2024; Wang et al., 2025; Edge et al., 2024). Then a subgraph is retrieved by expanding the local neighborhood of these seed nodes or relations. The structured subgraph and its associated content are incorporated into a prompt template to support LLM-based answer generation. However, as discussed in Section 1, existing methods exhibit two key limitations. First, the constructed KG is incomplete, which affects its effectiveness in supporting accurate answer generation. Second, they suffer from semantic misalignment due to limited query integration during retrieval and heavy reliance on LLMs for iterative query reformulation.
3. Our Proposed Approach: CUE-RAG
Our Goals. The objective of this work is to improve the effectiveness of graph-based RAG by addressing the limitations in existing methods. Specifically, we aim to: (1) design a multi-partite graph index that supports semantically coherent retrieval across multiple levels of granularity; (2) develop a hybrid strategy for knowledge unit extraction that balances LLM utility with computational cost; and (3) propose Q-Iter, an iterative retrieval strategy that improves retrieval accuracy.
We provide an overview of CUE-RAG in Section 3.1 and CUE-RAG’s architecture, comprising two core modules: CUE-Index for offline indexing in Section 3.2 and Q-Iter for online retrieval in Section 3.3.
@@ 3.1. Overview
At a high level, CUE-RAG consists of two modules: offline indexing CUE-Index and online retrieval Q-Iter. As shown in Figure 3, offline CUE-Index has two stages:
  1. (1)
    Multi-Partite Graph Construction: Algorithm 1 constructs a multi-partite graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ by linking chunk nodes to their corresponding knowledge unit nodes, and knowledge unit nodes to the entities they reference. This three-tier structure enables semantically coherent retrieval across multiple levels of granularity.
  2. (2)
    Hybrid Extraction: Given a text corpora $\mathcal{T}$ and a token constraint coefficient $\alpha\in[0,1]$, a subset of core text chunks $\mathcal{T}_{c}\in\mathcal{T}$ is selected based on relevance metrics. For $t\in\mathcal{T}_{c}$, LLM is employed to extract disambiguated knowledge units. For $t\in\mathcal{T}\setminus\mathcal{T}_{c}$, NLP tool segments the text into sentence-level units. Finally, entities are extracted from all resulting knowledge units using NLP tool.
During online retrieval, Q-Iter begins by extracting entities from the input query $q$ to activate related entities, while semantic search anchors initial knowledge units. These units expand the entity set via referenced entities, enabling constrained multihop traversal over $\mathcal{G}$ to retrieve additional relevant units. The resulting units are mapped to their source chunks and re-ranked to form the final context for answer generation.
3.2. Offline CUE-Index
Multi-Partite Graph Construction. CUE-Index constructs a multi-partite graph $\mathcal{G}$ from a text corpora $\mathcal{T}$, organizing information across three semantic levels, as illustrated in Figure 3. In the upper layer, chunk nodes $\mathcal{V}_{\mathcal{T}}$ represent coarse-grained content by encoding text chunks. The intermediate layer contains knowledge unit nodes $\mathcal{V}_{\mathcal{K}}$, which capture semantically richer information distilled from the text. In the lower layer, entity nodes $\mathcal{V}_{e}$ encode fine-grained factual elements extracted from knowledge units. The graph is then connected by two types of edges: $\mathcal{E}_{c}$, linking each chunk node to its corresponding knowledge units, and $\mathcal{E}_{e}$, connecting knowledge units to the entities they mention.
Hybrid Extraction. As shown in Algorithm 1, the hybrid extraction process begins by selecting a subset of core text chunks $\mathcal{T}_{c}\subseteq\mathcal{T}$ using Algorithm 2, based on a token constraint coefficient $\alpha\in[0,1]$. Algorithm 2 formulates the selection stage as a 0-1 knapsack problem, where each chunk is assigned a value based on its relevance score and a weight corresponding to its token length, selecting the subset that maximizes relevance under the token budget. In Algorithm 1, once $\mathcal{T}_{c}$ is identified (Line 3), each text chunk $t_{i}$ is processed according to its selection status (Lines 6–11). For $t_{i}\in\mathcal{T}_{c}$, a LLM extracts disambiguated knowledge units $\mathcal{K}_{i}$ that capture core semantics. For $t_{i}\in\mathcal{T}\setminus\mathcal{T}_{c}$, lightweight NLP tools (Bird et al., 2009) segment the text into sentences as knowledge units $\mathcal{K}_{i}$ to ensure broad coverage at lower computational cost. All text chunks and extracted knowledge units are instantiated as text nodes $\mathcal{V}_{\mathcal{T}}$ and knowledge unit nodes $\mathcal{V}_{\mathcal{K}}$, respectively, forming the upper and intermediate layers of the multi-partite graph. Finally, in Lines 12-13, the Named Entity Recognition (NER) component in spaCy (Honnibal et al., 2020) is utilized to extract entities $\mathcal{V}_{k}$ for each knowledge unit $k$. All these extracted entities $\mathcal{V}_{k}$ are collected and form entity nodes $\mathcal{V}_{e}$, acting as anchors for precise information retrieval.
@@ 3.2. Offline CUE-Index
3Output: A hierarchical graph index $\mathcal{G}=(\mathcal{V},\mathcal{E})$

### v3

# Clue-RAG: Towards Accurate and Cost-Efficient Graph-based RAG via Multi-Partite Graph and Query-Driven Iterative Retrieval

@@ Abstract.
Despite the remarkable progress of Large Language Models (LLMs), their performance in question answering (QA) remains limited by the lack of domain-specific and up-to-date knowledge. Retrieval-Augmented Generation (RAG) addresses this limitation by incorporating external information, often from graph-structured data. However, existing graph-based RAG methods suffer from poor graph quality due to incomplete extraction and insufficient utilization of query information during retrieval. To overcome these limitations, we propose Clue-RAG, a novel approach that introduces (1) a multi-partite graph index incorporates text Chunk, knowledge unit, and entity to capture semantic content at multiple levels of granularity, coupled with a hybrid extraction strategy that reduces LLM token usage while still producing accurate and disambiguated knowledge units, and (2) Q-Iter, a query-driven iterative retrieval strategy that enhances relevance through semantic search and constrained graph traversal. Experiments on three QA benchmarks show that Clue-RAG significantly outperforms state-of-the-art baselines, achieving up to 99.33% higher Accuracy and 113.51% higher F1 score while reducing indexing costs by 72.58%. Remarkably, Clue-RAG matches or outperforms baselines even without using an LLM for indexing. These results demonstrate the effectiveness and cost-efficiency of Clue-RAG in advancing graph-based RAG systems. The source code is available in https://github.com/Feesuu/ClueRAG.
@@ 1. Introduction
Large Language Models (LLMs) like Qwen3 (Yang et al., 2025), DeepSeek (DeepSeek-AI, 2025), and LLaMA3.1 (Grattafiori et al., 2024) have received tremendous attention from both industry and academia (Ghimire et al., 2024; Huang and Huang, 2024; Liu et al., 2024; Nie et al., 2024; Wang et al., 2024). Despite their remarkable success in question-answering (QA) tasks, they may still generate wrong answers due to a lack of domain-specific and real-time updated knowledge outside their pre-training corpus (Peng et al., 2024). To enhance the trustworthiness and interpretability of LLMs, Retrieval-Augmented Generation (RAG) methods (Fan et al., 2024; Gao et al., 2023; Hu and Lu, 2024; Huang and Huang, 2024; Wu et al., 2024a; Yu et al., 2024; Zhao et al., 2024; Wu et al., 2024b) have emerged as a core approach, which often retrieve relevant information from documents, relational data, and graph data to facilitate the QA tasks. The state-of-the-art (SOTA) approaches often use the graph to model the external data since they capture the rich semantic information and link relationships between entities.
@@
Figure 1. Comparison of triple extraction in existing Graph-based RAG methods (Upper) and ours (Below).
In the literature, representative graph-based RAG methods typically follow a two-phase pipeline: (1) Offline index construction: They first segment the external text corpora $\mathcal{T}$ into small chunks, then extract the nodes and edges, together with their associated textual attributes (e.g., descriptions (Edge et al., 2024) or keywords (Guo et al., 2024)) from chunks using LLMs, and finally build a knowledge graph (KG) (Zhou et al., 2025; Xiang et al., 2025). (2) Online retrieval: Given an online query $q$, they either retrieve relevant nodes, edges, or subgraphs from the KG, or optionally trace these elements back to their source text chunks via associated provenance links (Gutiérrez et al., 2024; Edge et al., 2024; Guo et al., 2024). Afterwards, they incorporate the relevant information from KG or chunks into a prompt template, and then feed the prompt into an LLM for answer generation. Although these methods offer improved performance, they still suffer from two major limitations.
$\bullet$ Limitation 1: the indexes suffer from the incompleteness issue. As shown in recent studies (Zhou et al., 2025; Xiang et al., 2025), the constructed KGs often suffer from incompleteness, typically due to missing key nodes, edges, or the omission of information that cannot be incorporated into structured graph elements. Take the text chunk in Figure 1 as an example. Existing methods typically extract only a few triples (e.g., $\langle\textit{Barcelona},\textit{won},\textit{Copa del Rey}$  $\textit{ first legagainst Getafe}\rangle$), while completely ignoring valuable textual information, such as nuanced comparisons between Messi’s performance and Maradona’s “Goal of the Century”. This incompleteness may lead to retrieval errors, i.e., when a query involving this detail is issued, the KG fails to provide sufficient information for generating an accurate answer, as shown in Figure 1.
$\bullet$ Limitation 2: the retrieval methods face the issue of semantic misalignment. In the online retrieval phase, most existing graph-based RAG methods (Huang et al., 2025; Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) follow two steps: (1) use the input query to identify relevant elements (e.g., nodes, edges) in the KG, and (2) retrieve additional information based on the above elements, such as text chunks or subgraphs, to augment the context for answer generation. While step (1) ensures query relevance, step (2) often fails to fully leverage the query information, leading to a semantic misalignment between the query and the retrieved context.
@@
To address the above limitations, this paper introduces Clue-RAG, a novel graph-based RAG approach, which not only builds a high quality graph, but also well utilize the query information for answering the questions. For Limitation 1, we introduce a novel concept, called knowledge unit, which is a statement that conveys an atomic piece of information from a text chunk. Collectively, these knowledge units can fully reconstruct the semantic content of the original text. By extracting entities from the knowledge units, we can effectively address missing nodes or edges in the constructed KG. As a result, we build a multi-partite graph composed of three types of nodes, each representing distinct semantic granularities: text Chunk, knowledge unit, and entity. These nodes are connected through extraction and containment relations, forming a coherent multi-partite graph, as illustrated in Section 3.2 and Figure 3.
@@
Intuitively, we can use some lightweight NLP tools (Bird et al., 2009) to extract sentences from a text chunk, and each sentence can serve as a knowledge unit. However, this approach overlooks sentence context, leading to contextual ambiguity – semantically similar sentences from different chunks may actually have distinct meanings. In Figure 2, for example, if we use an NLP tool to extract sentences from the two chunks, the results may have high cosine similarity (0.91) despite differing in meaning. To address this, we leverage LLMs to incorporate context when extracting knowledge units. As a result, the LLM-extracted units in the same example show much lower similarity (0.51), effectively capturing context-sensitive distinctions. However, directly using LLMs to extract the units for all text chunks is very time and token-consuming.
In this paper, we propose a novel hybrid extraction strategy that combines the advantages of powerful LLMs and lightweight NLP tools (Bird et al., 2009). An ideal strategy would involve an oracle model that can identify which knowledge unit in a chunk may exhibit contextual ambiguity with units in other chunks, and then use an LLM to resolve the ambiguous unit, while relying on NLP tools for unambiguous units. However, such an oracle model does not exist in practice. In light of this, we propose to use semantic similarity between chunks as a proxy to evaluate potential contextual ambiguity, since knowledge units in highly similar chunks are intuitively more prone to ambiguity. Specifically, we assign chunks having high semantic similarity with others for LLMs processing, while the remainder use NLP tools. To balance token cost and extraction accuracy, we formulate the chunk assignment as a 0-1 knapsack problem (Sahni, 1975), aiming to maximize the total semantic similarity of the chunks selected for LLM processing within a given token budget. As shown in our later experiments, this strategy matches the performance of end-to-end LLM-based unit extraction, but only uses 50% of the tokens. Remarkably, Clue-RAG matches or outperforms baselines even without using an LLM for indexing.
Figure 2. Comparing the knowledge units extracted by NLP tool and LLM.
@@
To address Limitation 2, we propose Query-driven iterative retrieval (Q-Iter), an iterative retrieval strategy that alternately traverses knowledge units and entities. The process begins by extracting entities from the query to activate related entities, while concurrently using semantic search to anchor associated knowledge units. These units, in turn, expand the entity set by introducing entities referenced in their content. The strategy then performs constrained multihop traversal over the multi-partite graph to retrieve more relevant units. To ensure alignment with the query, each candidate unit is scored using a lightweight re-ranker specialized in query-context relevance. Finally, the collected knowledge units are mapped to their originating text chunks, which are subsequently re-ranked to yield the final context for answer generation.
@@
In our experiments, we test 13 solutions across three datasets, focusing on two key aspects: (1) Cost efficiency, based on token usage during offline indexing and online retrieval; and (2) QA performance, measured by F1 score and Accuracy. Remarkably, Clue-RAG demonstrates significant improvements over the SOTA baselines, delivering up to 99.33% higher Accuracy and 113.51% greater F1 score while simultaneously reducing indexing costs by 72.58%. Besides, the zero-token variant of Clue-RAG achieves comparable or superior performance to baselines. This highlights Clue-RAG’s superior graph indexing and inherent retrieval capabilities.
In summary, we make the following principal contributions:
@@
  • We propose a novel multi-partite graph index that integrates text chunks, knowledge units, and entities to enable semantically coherent cross-granularity retrieval, coupled with a hybrid extraction strategy for knowledge units that strategically balances the benefits of LLM processing against token usage.
@@
  • We conduct extensive experiments and demonstrate the superior performance of Clue-RAG over baselines.
Outline. We provide the preliminaries and problem formulation in Section 2. Section 3.1 presents an overview of Clue-RAG, Section 3.2 details the construction of the offline index Clue-Index, and Section 3.3 describes the online retrieval algorithm Q-Iter. The experimental results are then reported in Section 4. We review related work in Section 5 and conclude the paper in Section 6.
Figure 3. Clue-RAG consists of two phases: offline indexing and online retrieval.
@@ 2. Preliminary
Let $\mathcal{T}$ be a text corpora whose element $t_{i}\in\mathcal{T}$ is a text chunk whose token length is at most $L$. The embedding of each text chunk $t_{i}$ is denoted by $\bm{\phi}(t_{i})$. We denote a graph by $\mathcal{G}=(\mathcal{V,E})$, where $\mathcal{V}$ represents the set of nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of edges connecting pairs of nodes. In addition, we use calligraphic uppercase letters (e.g., $\mathcal{S}$) to represent a set of elements. Table 1 provides a summary of frequently used notations.
@@
NN Returned chunks size NN
@@
Problem 1 (Graph-based RAG).
Given an LLM, a text corpora $\mathcal{T}$, and a set of questions $Q$, find the relevant information for $Q$ from the KG built on $\mathcal{T}$, which can be used to augment $Q$ so that the LLM can generate high-quality answers for $Q$.
3. Our Proposed Approach: Clue-RAG
In this section, we develop a novel graph-based RAG, called Clue-RAG, which addresses the two major limitations in existing methods by manipulating text Chunk, knowledge unit, and entity. Specifically, we design a multi-partite graph index that supports semantically coherent retrieval across multiple levels of granularity, coupled with a hybrid strategy for knowledge unit extraction that balances LLM utility with computational cost. Besides, we propose Q-Iter, an iterative retrieval strategy that improves retrieval accuracy. We first provide an overview of Clue-RAG in Section 3.1 and then describe its two core modules in Sections 3.2 and 3.3.
@@ 3.1. Overview
At a high level, Clue-RAG consists of two modules: offline indexing Clue-Index and online retrieval Q-Iter. As shown in Figure 3, the offline Clue-Index has two stages:
  • Stage-1: hybrid extraction. Given a text corpora $\mathcal{T}$ and a token constraint coefficient $\alpha\in[0,1]$, a subset of core text chunks $\mathcal{T}_{c}\subseteq\mathcal{T}$ is selected based on relevance metrics. For each $t\in\mathcal{T}_{c}$, an LLM is employed to extract disambiguated knowledge units. For $t\in\mathcal{T}\setminus\mathcal{T}_{c}$, we use an NLP tool to segment the text into sentence-level units. Finally, entities are extracted from knowledge units using an NLP tool.
  • Stage-2: graph construction. We construct a multi-partite graph index $\mathcal{G}=(\mathcal{V},\mathcal{E})$, including three types of nodes: chunk, knowledge unit, and entity. In $\mathcal{G}$, edges only connect nodes of different types: specifically, chunk nodes are linked to their corresponding knowledge unit nodes, and knowledge unit nodes are further connected to the entities they reference.
During the online retrieval, Q-Iter begins by extracting entities from the input query $q$ to activate related entities, while semantic search anchors initial knowledge units. These units expand the entity set via referenced entities, enabling constrained multihop traversal over $\mathcal{G}$ to retrieve additional relevant units. The resulting units are mapped to their source chunks and re-ranked to form the final context for answer generation.
@@
3Output: A multi-partite graph index $\mathcal{G}=(\mathcal{V},\mathcal{E})$